# GPT-Image-2 in the Wild: A Twitter Dataset of Self-Reported AI-Generated Images from the First Week of Deployment

### 28 Apr 2026 v1

@@ 3.2. Query Design and Collection
6 Query Types hashtags, @mentions, model name Twitter API v2 27,662 records Stage 1 Media Filter Stage 2 Text Heuristics Text-Confirmed 4,187 images Uncertain $\sim$21,500 Rejected $\sim$780 Provenance Validation Playwright / Chromium Badge-Confirmed 4,750 images No badge / unevaluable 1,147-1{,}14726,515
@@ 5.4. Semantic Clustering via CLIP
We embed all 10,217 images with CLIP ViT-L/14 [7], reduce to 2-D via UMAP [5] (cosine metric, $k=30$ neighbours, min-dist 0.05), and cluster with HDBSCAN [1] (min cluster size 15). The procedure yields 137 distinct semantic clusters (Figure 10); 33.2% of images are assigned to noise, reflecting the genuine visual diversity of the collection. The largest clusters contain 603, 220, and 175 images respectively, each corresponding to visually coherent aesthetic families (e.g., anime group portraits, typographic posters, photorealistic headshots). The language-coloured projection (Figure 10, right) shows partial but not complete language segregation: Japanese and Chinese creators overlap substantially in the anime cluster regions, while English creators are more evenly distributed across clusters.
@@ 6. Conclusion
We have presented the GPT-Image-2 Twitter Dataset: 10,217 confirmed GPT-image-2 images collected from Twitter/X within six days of the model’s public release, curated through a multilingual text-heuristic pipeline spanning English, Japanese, and Chinese creation-language signals, browser-automated Twitter badge verification, and model name variant matching. We document that C2PA content credentials are systematically stripped by Twitter’s CDN, and that Twitter’s native “Made with AI” badge (present on 53.7% of checked uncertain tweets) provides a complementary platform-level provenance signal. The dataset and all associated code are released at [REPOSITORYURL].

### 05 May 2026 v2

@@ 3.2. Query Design and Collection
6 Query Types hashtags, @mentions, model name Twitter API v2 27,662 records Stage 1 Media Filter Stage 2 Text Heuristics Text-Confirmed 5,467 images Uncertain $\sim$21,500 Rejected $\sim$780 Provenance Validation Playwright / Chromium Badge-Confirmed 4,750 images No badge / unevaluable 1,147-1{,}14726,515
@@ 5.4. Semantic Clustering via CLIP
We embed all 10,217 images with CLIP ViT-L/14 [6], reduce to 2-D via UMAP [4] (cosine metric, $k=30$ neighbours, min-dist 0.05), and cluster with HDBSCAN (min cluster size 15). The procedure yields 137 distinct semantic clusters (Figure 9); 33.2% of images are assigned to noise, reflecting the genuine visual diversity of the collection. The largest clusters contain 603, 220, and 175 images respectively, each corresponding to visually coherent aesthetic families (e.g., anime group portraits, typographic posters, photorealistic headshots). The language-coloured projection (Figure 9, right) shows partial but not complete language segregation: Japanese and Chinese creators overlap substantially in the anime cluster regions, while English creators are more evenly distributed across clusters.
@@
6. Limitations
A material caveat for downstream use is that GPT-image-2 supports both text-to-image generation (prompting from scratch) and image-to-image editing (modifying an existing photo). Our confirmation signals—creation-language phrases (“made with,” “generated by,” “prompt:”) and Twitter’s “Made with AI” badge—do not distinguish between the two: creators use the same vocabulary for both workflows, and Twitter applies the badge identically. The dataset is therefore a mixture of fully-generated and AI-edited images, and we do not have a reliable signal to separate them at scale because creators rarely disclose which workflow they used. This has implications for downstream research: detectors trained on this dataset will learn the signatures of both regimes together; attribution to GPT-image-2 specifically is harder for edited images where most pixels may be unmodified; and the face, OCR, and demographic statistics in Section 5 include features that may have been carried over from input photos rather than synthesised.
@@ 7. Conclusion
We have presented the GPT-Image-2 Twitter Dataset: 10,217 confirmed GPT-image-2 images collected from Twitter/X within six days of the model’s public release, curated through a multilingual text-heuristic pipeline spanning English, Japanese, and Chinese creation-language signals, browser-automated Twitter badge verification, and model name variant matching. We document that C2PA content credentials are systematically stripped by Twitter’s CDN, and that Twitter’s native “Made with AI” badge (present on 53.7% of checked uncertain tweets) provides a complementary platform-level provenance signal. The collection and curation code is released at https://github.com/scamai/gpt-image-2-dataset; the dataset itself is released separately at https://www.scam.ai/en/research.